OpenStudy (anonymous):

x^2-5x+5=0 what are the solutions to approx 3 decimal places

OpenStudy (lilg132):

use quadratic formula

OpenStudy (lilg132):

do you know the quadratic formula

OpenStudy (anonymous):

this is what ive gotten so far x-8= \[-\sqrt{20}\]

OpenStudy (anonymous):

\[x-8=-\sqrt{20}\]

OpenStudy (lilg132):

using the formula ? remember a = 1 b = -5 c = 5

OpenStudy (anonymous):

This one seems to only have one solution. Apply the following formula:\[\frac{-b \pm \sqrt{b^2 - 4 a c}}{2a}\]where a, b, and c correspond to the coefficients of your quadratic equation \[ax^2 + bx + c = 0\]

OpenStudy (lilg132):

and because b = -5 and it says to input -b in the formula just remember - * - = + so where it says -b you should put 5 not -5

OpenStudy (anonymous):

thats whats throwing me off is the abc formula

OpenStudy (lilg132):

what is

OpenStudy (anonymous):

when i have to put it into the different formula